# Characterisation of a quantum bus between two driven qubits

### v2

@@ I Introduction
In this work, we investigate the effects of dissipation and thermal effects on dressed qubits coupled to a harmonic oscillator. First, we analyze how the qubit’s state can be inferred by measuring a resonance frequency shift of the harmonic oscillator. In the dispersive regime, such a measurement corresponds to a quantum nondemolition measurement of the qubit. This mechanism has been demonstrated in transmission line resonators [19, 20, 21, 22, 23]. The effect of dissipation on the system due to its coupling to the environment, modeled as a heat bath, depends on the dressing of the qubit. Interestingly, we find that the qualitative form of dissipation differs from that of bare qubits: whereas in bare qubits the steady state is the ground state, corresponding to one of the qubit eigenstates, a dressed qubit relaxes towards a mixed state. The latter makes the two logical qubit states symmetric in terms of the unwanted dissipation. This may be advantageous since most quantum computing protocols do not differentiate the qualitative properties of the two logical states. We also demonstrate that the state readout mechanism is more robust for a dressed qubit compared to a bare one, particularly at higher temperatures. Following the approach of Refs. [19, 17, 24, 25], we study how the harmonic-oscillator-mediated coupling can be utilised to implement a two-qubit gate. We show how the thermal occupation of the harmonic oscillator affects the ability to generate qubit-qubit entanglement. Additionally, we compare gate fidelities across different model parameters and propose an optimal parameter range for practical implementation. Notably, we show that the fidelity exhibits a non-monotonic dependence on the dressing rise time. This non-monotonic behaviour can be leveraged to enhance fidelity in setups with a finite dressing speed.
@@
The work is organised as follows. In Sec. II, we introduce the quantum master equation formalism used to study the qubit-harmonic oscillator system. In Sec. III, we investigate how the damping and temperature affect the qubit state readout mechanism, comparing the performance of dressed and bare qubits and highlighting the advantages of using dressed qubits. In Sec. IV, we analyze how various system parameters influence the fidelity of a two-qubit gate mediated by a harmonic oscillator and propose an optimal parameter range for practical implementation. In Sec. V, we discuss possible realisations of the qubit-harmonic oscillator system across different technological platforms. Finally, we summarise the results in Sec. VI.
@@ II.1 Dressed qubit-harmonic oscillator coupling
where $\hat{\sigma}_{i}$ are the Pauli matrices describing the qubit degrees of freedom, $\hat{a}^{\dagger}$ and $\hat{a}$ are the raising and lowering operators of the harmonic oscillator, and $g$ is the coupling strength, which depends on the coupling mechanism and its implementation. As described in Supplementary note A [29], within the rotating wave approximation, the coupling term describes the exchange of energy quanta between the qubit and the harmonic oscillator. If $\omega_{q}$ is of the order of $\omega_{h}$, the coupling between the subsystems leads to a sizable interaction between them. However, the characteristic frequencies of the qubit and the bus may differ by several orders of magnitude. If the qubit transition frequency $\omega_{q}$ is much higher than $\omega_{h}$, the qubit may be driven by an input field in resonance with $\omega_{q}$ to dress the qubit with the desired eigenfrequency of the Rabi oscillations. For instance, a spin qubit can be dressed by applying a microwave field. Its magnetic field component, pointing in a direction perpendicular to the dc field, dresses the spin qubit, as shown in Fig. 1(a). Here, we treat the driving field as a classical field interacting with the qubit. The Hamiltonian describing the dressed qubit-harmonic oscillator system then reads:
@@
where the first two terms in the Hamiltonian describe the dressed qubit. For a drive in resonance with the bare qubit $\Delta=0$, and within the rotating wave approximation (note that this is distinct from the the rotating wave approximation mentioned for the qubit-harmonic oscillator coupling), Eq. (3) becomes the well-known Jaynes–Cummings Hamiltonian, see Supplementary note A [29]. The eigenstates of $\hat{\sigma}_{x}$ describe the logical states $\{|0\rangle,|1\rangle\}$ of the qubit. The qubit can be selectively coupled to the harmonic oscillator by changing the driving amplitude $\Omega_{R}(t)$. In Sec. IV.2, we study how the qubit coupling speed affects the fidelity of a harmonic oscillator-mediated gate.
As shown in Supplementary note A [29], if the dressed qubit is near-resonant with the harmonic oscillator, $\Omega_{R}\sim\omega_{h}$, the coupling between them shifts the resonance frequency of the oscillator in a way that depends on the qubit state. Therefore, the state of the qubit can be detected by measuring the noise power spectral density of the $\hat{x}(t)=\hat{a}+\hat{a}^{\dagger}$ quadrature
@@ II.2 Quantum master equation
In Eq. (7) we ignore the renormalisation of the system Hamiltonian introduced by the system-bath coupling, since the bare parameters are not accessible. $\hat{L}_{k}$ are Lindblad operators describing the effect of the baths on the system. They are related to the system operators coupling to the baths as
@@
where $n_{\mathrm{th},q}=(e^{\omega_{q}/T}-1)^{-1}$ and $n_{\mathrm{th},h}=(e^{\omega_{h}/T}-1)^{-1}$ describe the thermal state of the bath fields coupling to the qubit and the harmonic oscillator, respectively. For the dressed qubit, we assume that the bare transition frequency is much higher than the temperature, i.e., $\omega_{q}\gg T$, so that $n_{\mathrm{th},q}=0$. The terms containing the Lindblad operators describing the coupling of the harmonic oscillator to the heat bath, $\hat{L}_{h,\pm}$, induce damping, driving the harmonic oscillator towards a thermal equilibrium state with the heat bath. Similarly, the terms containing $\hat{L}_{q,\pm}$ thermalise a bare qubit with its environment. Here, we assume that the qubit damping is stronger than pure dephasing, and only consider the Lindblad operators $\hat{L}_{q,\pm}$ describing the damping.
@@
III Characterisation of the qubit-harmonic oscillator coupling
@@
In the dispersive regime $|\Delta_{R}|\gg g$, the resonance frequency of the harmonic oscillator gets shifted by ±$\pm\frac{g^{2}}{\Delta_{R}}$ depending on the qubit’s state [19, 22], see Eq. (A.2) in the Supplementary Information [29]. The resonance frequency manifests as a maximum in the noise power spectral density, Eq. (4). The shift of the maximum can be detected, for instance, in an optomechanical quantum bus, through optical measurements, and pulsed measurements allow for the detection of short-lived shifts in the mechanical frequency [28, 33, 34, 16]. The resonance frequency shift is generally measurable if it is greater than the linewidth.
In this section, we analyze the frequency- and time-domain response of a qubit coupled to a harmonic oscillator. We solve the QME (7) for the system Hamiltonian (3) [dressed qubit] and (1) [bare qubit]. The codes used to solve the equation are available at [35]. The temperature $T$ of the heat bath determines the thermal occupations $n_{\mathrm{th},q}$ and $n_{\mathrm{th},h}$ in the Lindblad operators, and the initial state of the harmonic oscillator, which we assume to be in thermal equilibrium with the heat bath so that $\rho_{h}(0)=Z^{-1}e^{-\omega_{h}\hat{a}^{\dagger}\hat{a}/T}$, where $Z$ is the partition function of the harmonic oscillator. The qubit is initialised either in the ground ($|0\rangle$) or the excited ($|1\rangle$) state. The noise power spectral density introduced in Eq. (4) is written in the Heisenberg picture, while the QME formalism is based on the Schrödinger picture. Using the quantum regression theorem [36, 37], the two-time correlation function in Eq. (4) can be rewritten as
@@
As shown in Fig. 2(a), in the absence of thermal occupation of the oscillator (blue line), the resonance frequency shift for the dressed qubit is given approximately by ±$\pm\frac{g^{2}}{\Delta_{R}}$ [Eq. (A.2) in the Supplementary Information [29]], with the upper and lower signs corresponding to the excited and ground states, respectively. The resonance frequency shift is slightly suppressed at higher temperatures, but the resonance peak height increases, making it easier to detect the resonance frequency. $S(\omega)$ shows a small kink at the frequency corresponding to the opposite state $\mp\frac{g^{2}}{\Delta_{R}}$. This feature can be explained by studying the time-domain response. In Fig. 2(c) we show the time evolution of a qubit with initial state $|1\rangle$. The system begins at the excited state with $\langle\sigma_{x}\rangle=1$ (see Hamiltonian (3)), and evolves into a mixed state due to the damping [31]. As shown in Supplementary note B [29], in the steady state, the weight of the ground state is higher than that of the excited state, so that $\langle\sigma_{x}\rangle<0$. The dashed line in Fig. 2(c) indicates the value of $\langle\sigma_{x}\rangle$ in the steady state, given by Eq. (B.4) in the Supplementary Information [29]. The contribution of the $|0\rangle$ state to the noise power spectral density leads to the kinks at $\omega=\omega_{h}-\frac{g^{2}}{\Delta_{R}}$.
@@ IV Implementation of a two-qubit gate
In the previous section, we characterise the coupling between the qubit and the harmonic oscillator, highlighting the benefits of using a driven qubit for state readout. In the following, we consider two driven qubits coupled to the same harmonic oscillator, as illustrated in Fig. 1(b). We study the implementation of a two-qubit gate, where the qubit-qubit interaction is mediated by the harmonic oscillator. The Hamiltonian for a two-qubit system can be obtained by a direct generalisation of Hamiltonian (3)
@@
Within the rotating wave approximation, the coupling between the qubit and the harmonic oscillator gives rise to an effective qubit-qubit interaction [38, 19, 24], see Supplementary note C [29]. The interaction time between the qubits can be controlled by tuning their drive amplitudes $\Omega_{R,j}(t)$ close to resonance with the harmonic oscillator. Assuming equal drive amplitudes and coupling strengths for both qubits, the unitary evolution of the system for an interaction time of $t_{\mathrm{int}}=\frac{\pi|\Delta_{R}|}{4g^{2}}$ corresponds to a $\sqrt{i\mathrm{SWAP}}$-gate operation, up to one-qubit transformations,
@@
where the upper and lower signs correspond to a positive and negative Rabi detuning. The $\sqrt{i\mathrm{SWAP}}$-gate generates entanglement between the qubits if they are in opposite logical states. For instance, it maps the separable state $|10\rangle$ into the maximally entangled state $(|10\rangle+i|01\rangle)/\sqrt{2}$. In Supplementary note D [29] we show how to create Bell states in driven qubits. In the following, we study how the qubit-harmonic oscillator coupling enables the qubit-qubit entanglement generation, and analyze the gate fidelity to determine a suitable range of the system parameters for a practical implementation.
@@ IV.1 Qubit entanglement
Universal quantum computation requires the capacity to generate any pure state from an arbitrary initial state. Typically, qubits are initialised in a separable state, which can be manipulated into an entangled state through a sequence of quantum gates. In this section, we analyze the capacity to generate a fully entangled two-qubit state. The qubit-qubit coupling is mediated by a harmonic oscillator, which may become entangled with the qubits during the process. The main goal of this section is to determine whether the qubit-harmonic oscillator system results in qubit-qubit entanglement, or if these are independent processes, and to determine how thermal effects hinder the generation of a fully entangled state. Entanglement in pure bipartite systems is typically measured by the von Neumann entropy of the reduced density matrices. However, for mixed or multipartite systems, the von Neumann entropy is not always a suitable measure, as a classically correlated mixed state can have nonzero von Neumann entropy even though it is not entangled. Several alternative measures of entropy exist for mixed systems [39, 40]. Here, we use logarithmic negativity [41, 42] as a measure of entanglement. Given two subsystems, $A$ and $B$, logarithmic negativity is defined as
@@
where $\hat{\rho}^{\mathrm{T}_{A}}$ is the partial transpose with respect to subsystem $A$ and $||\hat{X}||=\mathrm{Tr}\sqrt{\hat{X}^{\dagger}\hat{X}}$ is the trace norm. Unlike the entropy of entanglement, logarithmic negativity is entanglement monotone [43, 42] both for pure and mixed states. For 2-level systems such as qubits, logarithmic negativity is normalised such that $E_{\mathcal{N}}=1$ for a maximally entangled state.
We solve the quantum master equation (7) for the system Hamiltonian (10[35]. To compute the entanglement between the different subsystems, we perform partial traces over $\hat{\rho}(t)$ to obtain the relevant reduced density matrix. In Fig. 3(a), we show the evolution of the two-qubit states for two consecutive applications ($0\leq t\leq 2t_{\mathrm{int}}$) of the $\sqrt{-i\mathrm{SWAP}}$ gate, with initial state $|10\rangle$ and thermal occupation $n_{\mathrm{th},h}=0$. Ideally, the system should evolve into the $(|10\rangle-i|01\rangle)/\sqrt{2}$ state at $t=t_{\mathrm{int}}$, and to the $-i|01\rangle$ state at $t=2t_{\mathrm{int}}$, swapping the states of the qubits up to a phase factor. On top of the swapping of the qubit occupation numbers $\sigma_{x,i}$, there are small oscillations with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$ and an amplitude of the order of $2g^{2}/\Delta_{R}^{2}$. The cross-correlation $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ oscillates close to -1, since the qubits ideally remain in opposite states. The small oscillations arise due to deviation of the parameters from the dispersive regime, i.e., by retaining higher-order terms in the $g\ll|\Delta_{R}|$ expansion carried out in Supplementary note C [29]. In the plot, we have chosen values of $g$ and $\Delta_{R}$ satisfying $(\Delta_{R}+4g^{2}/\Delta_{R})t_{\mathrm{int}}\in 2\pi\mathbb{N}$, so that $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ is minimal at $t=t_{\mathrm{int}}$, thus maximizing the gate fidelity $\mathcal{F}$. Nonetheless, the crossing point of the occupation numbers $\langle\sigma_{x}^{1}\rangle=\langle\sigma_{x}^{2}\rangle=0$ occurs at a time slightly higher than $t=t_{\mathrm{int}}$. A finer tuning of $\Delta_{R}$ and $g$ could simultaneously optimise the qubit expectation values to enhance the fidelity.
@@
In Fig. 3(c-d) we show the entanglement, quantified by the logarithmic negativity (12), between the different subsystems, namely the harmonic oscillator ($h$), the subsystem formed by the two qubits ($q$), and each qubit ($q_{1}$ and $q_{2}$). The system begins in a separable state and evolves into a state with maximally entangled qubits at $t=t_{\mathrm{int}}$ (purple line). The harmonic oscillator-qubit entanglement $E_{h,q}$ (yellow line) oscillates periodically with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$, while the entanglement of the harmonic oscillator with each of the individual qubits oscillates in antiphase. The amplitudes of the oscillations of $E_{h,q_{1}}$ (blue line) and $E_{h,q_{2}}$ (red line) are modulated by their respective qubit’s state. Since the qubits are in near resonance with the harmonic oscillator, one may simplify the coupling term in Eq. (10) by applying the rotating wave approximation. Within the rotating wave approximation, doubly rotating terms are disregarded, so that the remaining terms describe energy quanta exchange between the qubit and the harmonic oscillator ($\hat{\sigma}_{-,x}^{j}\hat{a}^{\dagger}$ and $\hat{\sigma}_{+,x}^{j}\hat{a})$. Here, $\hat{\sigma}_{\pm,x}$ are the rising and lowering operators in the dressed qubit’s basis. Since at $T=0$ the harmonic oscillator is initialised in the ground state, at $t=0$ it may only couple to the first qubit, which starts in the excited state $|1\rangle$. The situation is reversed at $t=2t_{\mathrm{int}}$.

### v3

# Characterization of a quantum bus between two driven qubits

@@ I Introduction
In this work, we investigate the effects of dissipation and thermal effects on dressed qubits coupled to a harmonic oscillator. First, we analyze how the qubit’s state can be inferred by measuring a resonance frequency shift of the harmonic oscillator. In the dispersive regime, such a measurement corresponds to a quantum nondemolition measurement of the qubit. This mechanism has been demonstrated in transmission line resonators [19, 20, 21, 22, 23]. The effect of dissipation on the system due to its coupling to the environment, modeled as a heat bath, depends on the dressing of the qubit. Interestingly, we find that the qualitative form of dissipation differs from that of bare qubits: whereas in bare qubits the steady state is the ground state, corresponding to one of the qubit eigenstates, a dressed qubit relaxes towards a mixed state. The latter makes the two logical qubit states symmetric in terms of the unwanted dissipation. This may be advantageous since most quantum computing protocols do not differentiate the qualitative properties of the two logical states. We also demonstrate that the state readout mechanism is more robust for a dressed qubit compared to a bare one, particularly at higher temperatures. Following the approach of Refs. [19, 17, 24, 25], we study how the harmonic-oscillator-mediated coupling can be utilized to implement a two-qubit gate. We show how the thermal occupation of the harmonic oscillator affects the ability to generate qubit-qubit entanglement. Additionally, we compare gate fidelities across different model parameters and propose an optimal parameter range for practical implementation. Notably, we show that the fidelity exhibits a non-monotonic dependence on the dressing rise time. This non-monotonic behavior can be leveraged to enhance fidelity in setups with a finite dressing speed.
@@
The work is organized as follows. In Sec. II, we introduce the quantum master equation formalism used to study the qubit-harmonic oscillator system. In Sec. III, we investigate how the damping and temperature affect the qubit state readout mechanism, comparing the performance of dressed and bare qubits and highlighting the advantages of using dressed qubits. In Sec. IV, we analyze how various system parameters influence the fidelity of a two-qubit gate mediated by a harmonic oscillator and propose an optimal parameter range for practical implementation. In Sec. V, we discuss possible realizations of the qubit-harmonic oscillator system across different technological platforms. Finally, we summarize the results in Sec. VI.
@@ II.1 Dressed qubit-harmonic oscillator coupling
where $\hat{\sigma}_{i}$ are the Pauli matrices describing the qubit degrees of freedom, $\hat{a}^{\dagger}$ and $\hat{a}$ are the raising and lowering operators of the harmonic oscillator, and $g$ is the coupling strength, which depends on the coupling mechanism and its implementation. As described in Appendix A, within the rotating wave approximation, the coupling term describes the exchange of energy quanta between the qubit and the harmonic oscillator. If $\omega_{q}$ is of the order of $\omega_{h}$, the coupling between the subsystems leads to a sizable interaction between them. However, the characteristic frequencies of the qubit and the bus may differ by several orders of magnitude. If the qubit transition frequency $\omega_{q}$ is much higher than $\omega_{h}$, the qubit may be driven by an input field in resonance with $\omega_{q}$ to dress the qubit with the desired eigenfrequency of the Rabi oscillations. For instance, a spin qubit can be dressed by applying a microwave field. Its magnetic field component, pointing in a direction perpendicular to the dc field, dresses the spin qubit, as shown in Fig. 1(a). Here, we treat the driving field as a classical field interacting with the qubit. The Hamiltonian describing the dressed qubit-harmonic oscillator system then reads:
@@
where the first two terms in the Hamiltonian describe the dressed qubit. For a drive in resonance with the bare qubit $\Delta=0$, and within the rotating wave approximation (note that this is distinct from the the rotating wave approximation mentioned for the qubit-harmonic oscillator coupling), Eq. (3) becomes the well-known Jaynes–Cummings Hamiltonian, see Appendix A. The eigenstates of $\hat{\sigma}_{x}$ describe the logical states $\{|0\rangle,|1\rangle\}$ of the qubit. The qubit can be selectively coupled to the harmonic oscillator by changing the driving amplitude $\Omega_{R}(t)$. In Sec. IV.2, we study how the qubit coupling speed affects the fidelity of a harmonic oscillator-mediated gate.
As shown in Appendix A, if the dressed qubit is near-resonant with the harmonic oscillator, $\Omega_{R}\sim\omega_{h}$, the coupling between them shifts the resonance frequency of the oscillator in a way that depends on the qubit state. Therefore, the state of the qubit can be detected by measuring the noise power spectral density of the $\hat{x}(t)=\hat{a}+\hat{a}^{\dagger}$ quadrature
@@ II.2 Quantum master equation
In Eq. (7) we ignore the renormalization of the system Hamiltonian introduced by the system-bath coupling, since the bare parameters are not accessible. $\hat{L}_{k}$ are Lindblad operators describing the effect of the baths on the system. They are related to the system operators coupling to the baths as
@@
where $n_{\mathrm{th},q}=(e^{\omega_{q}/T}-1)^{-1}$ and $n_{\mathrm{th},h}=(e^{\omega_{h}/T}-1)^{-1}$ describe the thermal state of the bath fields coupling to the qubit and the harmonic oscillator, respectively. For the dressed qubit, we assume that the bare transition frequency is much higher than the temperature, i.e., $\omega_{q}\gg T$, so that $n_{\mathrm{th},q}=0$. The terms containing the Lindblad operators describing the coupling of the harmonic oscillator to the heat bath, $\hat{L}_{h,\pm}$, induce damping, driving the harmonic oscillator towards a thermal equilibrium state with the heat bath. Similarly, the terms containing $\hat{L}_{q,\pm}$ thermalize a bare qubit with its environment. Here, we assume that the qubit damping is stronger than pure dephasing, and only consider the Lindblad operators $\hat{L}_{q,\pm}$ describing the damping.
@@
III Characterization of the qubit-harmonic oscillator coupling
@@
In the dispersive regime $|\Delta_{R}|\gg g$, the resonance frequency of the harmonic oscillator gets shifted by ±$\pm\frac{g^{2}}{\Delta_{R}}$ depending on the qubit’s state [19, 22], see Eq. (A.2). The resonance frequency manifests as a maximum in the noise power spectral density, Eq. (4). The shift of the maximum can be detected, for instance, in an optomechanical quantum bus, through optical measurements, and pulsed measurements allow for the detection of short-lived shifts in the mechanical frequency [28, 32, 33, 16]. The resonance frequency shift is generally measurable if it is greater than the linewidth.
In this section, we analyze the frequency- and time-domain response of a qubit coupled to a harmonic oscillator. We solve the QME (7) for the system Hamiltonian (3) [dressed qubit] and (1) [bare qubit]. The codes used to solve the equation are available at [34]. The temperature $T$ of the heat bath determines the thermal occupations $n_{\mathrm{th},q}$ and $n_{\mathrm{th},h}$ in the Lindblad operators, and the initial state of the harmonic oscillator, which we assume to be in thermal equilibrium with the heat bath so that $\rho_{h}(0)=Z^{-1}e^{-\omega_{h}\hat{a}^{\dagger}\hat{a}/T}$, where $Z$ is the partition function of the harmonic oscillator. The qubit is initialized either in the ground ($|0\rangle$) or the excited ($|1\rangle$) state. The noise power spectral density introduced in Eq. (4) is written in the Heisenberg picture, while the QME formalism is based on the Schrödinger picture. Using the quantum regression theorem [35, 36], the two-time correlation function in Eq. (4) can be rewritten as
@@
As shown in Fig. 2(a), in the absence of thermal occupation of the oscillator (blue line), the resonance frequency shift for the dressed qubit is given approximately by ±$\pm\frac{g^{2}}{\Delta_{R}}$ [Eq. (A.2)], with the upper and lower signs corresponding to the excited and ground states, respectively. The resonance frequency shift is slightly suppressed at higher temperatures, but the resonance peak height increases, making it easier to detect the resonance frequency. $S(\omega)$ shows a small kink at the frequency corresponding to the opposite state $\mp\frac{g^{2}}{\Delta_{R}}$. This feature can be explained by studying the time-domain response. In Fig. 2(c) we show the time evolution of a qubit with initial state $|1\rangle$. The system begins at the excited state with $\langle\sigma_{x}\rangle=1$ (see Hamiltonian (3)), and evolves into a mixed state due to the damping [30]. As shown in Appendix B, in the steady state, the weight of the ground state is higher than that of the excited state, so that $\langle\sigma_{x}\rangle<0$. The dashed line in Fig. 2(c) indicates the value of $\langle\sigma_{x}\rangle$ in the steady state, given by Eq. (B.4). The contribution of the $|0\rangle$ state to the noise power spectral density leads to the kinks at $\omega=\omega_{h}-\frac{g^{2}}{\Delta_{R}}$.
@@ IV Implementation of a two-qubit gate
In the previous section, we characterize the coupling between the qubit and the harmonic oscillator, highlighting the benefits of using a driven qubit for state readout. In the following, we consider two driven qubits coupled to the same harmonic oscillator, as illustrated in Fig. 1(b). We study the implementation of a two-qubit gate, where the qubit-qubit interaction is mediated by the harmonic oscillator. The Hamiltonian for a two-qubit system can be obtained by a direct generalization of Hamiltonian (3)
@@
Within the rotating wave approximation, the coupling between the qubit and the harmonic oscillator gives rise to an effective qubit-qubit interaction [37, 19, 24], see Appendix C. The interaction time between the qubits can be controlled by tuning their drive amplitudes $\Omega_{R,j}(t)$ close to resonance with the harmonic oscillator. Assuming equal drive amplitudes and coupling strengths for both qubits, the unitary evolution of the system for an interaction time of $t_{\mathrm{int}}=\frac{\pi|\Delta_{R}|}{4g^{2}}$ corresponds to a $\sqrt{i\mathrm{SWAP}}$-gate operation, up to one-qubit transformations,
@@
where the upper and lower signs correspond to a positive and negative Rabi detuning. The $\sqrt{i\mathrm{SWAP}}$-gate generates entanglement between the qubits if they are in opposite logical states. For instance, it maps the separable state $|10\rangle$ into the maximally entangled state $(|10\rangle+i|01\rangle)/\sqrt{2}$. In Appendix D we show how to create Bell states in driven qubits. In the following, we study how the qubit-harmonic oscillator coupling enables the qubit-qubit entanglement generation, and analyze the gate fidelity to determine a suitable range of the system parameters for a practical implementation.
@@ IV.1 Qubit entanglement
Universal quantum computation requires the capacity to generate any pure state from an arbitrary initial state. Typically, qubits are initialized in a separable state, which can be manipulated into an entangled state through a sequence of quantum gates. In this section, we analyze the capacity to generate a fully entangled two-qubit state. The qubit-qubit coupling is mediated by a harmonic oscillator, which may become entangled with the qubits during the process. The main goal of this section is to determine whether the qubit-harmonic oscillator system results in qubit-qubit entanglement, or if these are independent processes, and to determine how thermal effects hinder the generation of a fully entangled state. Entanglement in pure bipartite systems is typically measured by the von Neumann entropy of the reduced density matrices. However, for mixed or multipartite systems, the von Neumann entropy is not always a suitable measure, as a classically correlated mixed state can have nonzero von Neumann entropy even though it is not entangled. Several alternative measures of entropy exist for mixed systems [38, 39]. Here, we use logarithmic negativity [40, 41] as a measure of entanglement. Given two subsystems, $A$ and $B$, logarithmic negativity is defined as
@@
where $\hat{\rho}^{\mathrm{T}_{A}}$ is the partial transpose with respect to subsystem $A$ and $||\hat{X}||=\mathrm{Tr}\sqrt{\hat{X}^{\dagger}\hat{X}}$ is the trace norm. Unlike the entropy of entanglement, logarithmic negativity is entanglement monotone [42, 41] both for pure and mixed states. For 2-level systems such as qubits, logarithmic negativity is normalized such that $E_{\mathcal{N}}=1$ for a maximally entangled state.
We solve the quantum master equation (7) for the system Hamiltonian (10[34]. To compute the entanglement between the different subsystems, we perform partial traces over $\hat{\rho}(t)$ to obtain the relevant reduced density matrix. In Fig. 3(a), we show the evolution of the two-qubit states for two consecutive applications ($0\leq t\leq 2t_{\mathrm{int}}$) of the $\sqrt{-i\mathrm{SWAP}}$ gate, with initial state $|10\rangle$ and thermal occupation $n_{\mathrm{th},h}=0$. Ideally, the system should evolve into the $(|10\rangle-i|01\rangle)/\sqrt{2}$ state at $t=t_{\mathrm{int}}$, and to the $-i|01\rangle$ state at $t=2t_{\mathrm{int}}$, swapping the states of the qubits up to a phase factor. On top of the swapping of the qubit occupation numbers $\sigma_{x,i}$, there are small oscillations with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$ and an amplitude of the order of $2g^{2}/\Delta_{R}^{2}$. The cross-correlation $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ oscillates close to -1, since the qubits ideally remain in opposite states. The small oscillations arise due to deviation of the parameters from the dispersive regime, i.e., by retaining higher-order terms in the $g\ll|\Delta_{R}|$ expansion carried out in Appendix C. In the plot, we have chosen values of $g$ and $\Delta_{R}$ satisfying $(\Delta_{R}+4g^{2}/\Delta_{R})t_{\mathrm{int}}\in 2\pi\mathbb{N}$, so that $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ is minimal at $t=t_{\mathrm{int}}$, thus maximizing the gate fidelity $\mathcal{F}$. Nonetheless, the crossing point of the occupation numbers $\langle\sigma_{x}^{1}\rangle=\langle\sigma_{x}^{2}\rangle=0$ occurs at a time slightly higher than $t=t_{\mathrm{int}}$. A finer tuning of $\Delta_{R}$ and $g$ could simultaneously optimize the qubit expectation values to enhance the fidelity.
@@
In Fig. 3(c-d) we show the entanglement, quantified by the logarithmic negativity (12), between the different subsystems, namely the harmonic oscillator ($h$), the subsystem formed by the two qubits ($q$), and each qubit ($q_{1}$ and $q_{2}$). The system begins in a separable state and evolves into a state with maximally entangled qubits at $t=t_{\mathrm{int}}$ (purple line). The harmonic oscillator-qubit entanglement $E_{h,q}$ (yellow line) oscillates periodically with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$, while the entanglement of the harmonic oscillator with each of the individual qubits oscillates in antiphase. The amplitudes of the oscillations of $E_{h,q_{1}}$ (blue line) and $E_{h,q_{2}}$ (red line) are modulated by their respective qubit’s state. Since the qubits are in near resonance with the harmonic oscillator, one may simplify the coupling term in Eq. (10) by applying the rotating wave approximation. Within the rotating wave approximation, doubly rotating terms are disregarded, so that the remaining terms describe energy quanta exchange between the qubit and the harmonic oscillator ($\hat{\sigma}_{-,x}^{j}\hat{a}^{\dagger}$ and $\hat{\sigma}_{+,x}^{j}\hat{a})$. Here, $\hat{\sigma}_{\pm,x}$ are the rising and lowering operators in the dressed qubit’s basis. Since at $T=0$ the harmonic oscillator is initialized in the ground state, at $t=0$ it may only couple to the first qubit, which starts in the excited state $|1\rangle$. The situation is reversed at $t=2t_{\mathrm{int}}$.